OpenStudy (anonymous):

Solve: 3(n-2)-2(3-n) = 4(n-3)

OpenStudy (aroub):

1. Distribute 2.Combine like terms 3. Find 'n' 3(n-2)-2(3-n) = 4(n-3) => 3n-6-6+2n=4n-12 => 3n-12+2n=4n-12 =>3n+2n-4n=12-12 =>n=0

OpenStudy (anonymous):

Thanks. I didn't know if I had it right or not.

OpenStudy (anonymous):

n = 0